# Towards Automatic Evaluation and High-Quality Pseudo-Parallel Dataset Construction for Audio Editing: A Human-in-the-Loop Method

### v1

@@ Introduction
Table 1: Subjective evaluation result. Detailed distribution of difference editing types is presented in AppendixH I.
Editing
Model or Method
Generative
TTA Backbone
Timesteps
System
ID
Mean Scores
Quality
Relevance
Faithfulness
SDEdit Tango2 75 sys1 3.621 2.896 3.699
125 sys2 3.622 3.420 3.375
175 sys3 3.548 3.503 2.893
Auffusion 75 sys4 3.602 3.268 3.426
125 sys5 3.487 3.461 2.885
175 sys6 3.492 3.531 2.742
DDIM Inversion Tango2 200 sys7 3.420 2.791 3.489
Auffusion 200 sys8 3.867 3.447 3.441
Edit-friendly
DDPM Inversion
Tango2 75 sys9 3.658 3.330 3.641
125 sys10 3.684 3.354 3.479
175 sys11 3.655 3.593 3.408
Auffusion 75 sys12 3.639 3.029 3.483
125 sys13 3.687 3.431 3.417
175 sys14 3.669 3.432 3.167
CycleDiffusion
DDPM Inversion
Tango2 75 sys15 3.592 2.848 3.595
125 sys16 3.445 3.184 3.313
175 sys17 3.293 3.146 3.054
Auffusion 75 sys18 3.463 2.922 3.423
125 sys19 3.360 3.103 3.047
175 sys20 3.262 3.159 2.893
AP-adapter Audioldm2 - sys21 3.353 2.857 2.653
Musicgen - - sys22 3.208 2.968 2.693
Human Edit - - sys23 3.883 3.878 3.988
@@
AuditScore is a human-annotated dataset comprising 6,360 pairs of original and edited audio samples, totaling 35.3 hours in duration. These samples are generated using 23 distinct system configurations across seven representative audio editing approaches: SDEdit (Meng et al. ), DDIM (Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2020), Edit-Friendly DDPM (Huberman-Spiegelglas, Kulikov, and Michaeli 2024; Manor and Michaeli 2024), CycleDiffusion (Wu and De la Torre 2023), AP-adapter (Tsai et al. 2024), MusicGen (Copet et al. 2023), and Human Edit(manual editing). For diffusion-based methods (SDEdit, DDIM, Edit-Friendly DDPM, and CycleDiffusion), we adopt two state-of-the-art generative backbones—Tango2 (Majumder et al. 2024) and Auffusion (Xue et al. 2024)—to synthesize the edited audio. To further promote diversity and enhance model generalization, each configuration is instantiated with three inversion timesteps (75, 125, and 175 out of 200), allowing finer control over the degree of editing. This setup is specifically applied to SDEdit, Edit-Friendly DDPM, and CycleDiffusion to explore a wider range of editing intensities and semantic variations.
@@
To construct a representative and diverse dataset, we began by selecting original audio samples and corresponding captions from the AudioCaps dataset (Kim et al. 2019). Captions were clustered into 100 semantic groups using GTE-Large text embeddings (Li et al. 2023), from which 2–3 audio samples were manually selected per cluster. For each selected audio, we crafted three types of editing instructions—Addition, Deletion, and Replacement—yielding a shared pool of 240 instructions across all systems. Additionally, several representative systems were assigned 30 unique instructions to further increase variation in editing scenarios. In total, AuditScore covers over 200 distinct audio event types, with editing operations introducing more than 150 novel events. The rich variety of methods, model architectures, editing parameter settings, and content diversity enables AuditScore to capture a wide spectrum of editing behaviors, audio qualities, and modeling challenges—making it a comprehensive and robust benchmark for training and evaluating audio editing evaluation models.
@@
To assess the edited audio samples generated by these 23 systems, we conducted a large-scale subjective listening study involving five domain experts with extensive experience in audio dataset curation and perceptual evaluation. Each expert independently rated the editing audio pairs along three key perceptual dimensions: Quality, Relevance, and Faithfulness. The final rating for each system was computed as the average of expert scores across all three dimensions. A comprehensive summary of system IDs and their evaluation results is presented in Table 1.
(a) Quality (Distribution)
(b) Relevance (Distribution)
(c) Faithfulness (Distribution)
(d) Quality (Heatmap)
(e) Relevance (Heatmap)
(f) Faithfulness (Heatmap)
Figure 1: Comparison of quality, relevance, and faithfulness across all systems and annotation types. Top: score distribution across samples. Bottom: system-level heatmaps from AuditScore.
Scoring Protocol and Analysis
@@
To comprehensively evaluate audio editing performance, we adopt a three-dimensional scoring protocol encompassing Quality, Relevance, and Faithfulness. These dimensions are informed by established perceptual evaluation methodologies—such as MOS (Mean Opinion Score), CMOS (Comparison MOS)(Loizou 2011), and semantic alignment assessments—while tailored specifically for the nuances of audio editing. Detailed annotation guidelines were developed to ensure consistency across expert raters while preserving flexibility for nuanced perceptual judgments. The scoring scales are defined as follows(Additional scoring rubrics and details can be found in AppendixA B):
@@
Figure 1a–c show the distribution of scores across the three dimensions. The scores for Quality, Relevance, and Faithfulness each approximately follow a normal distribution, indicating that our scoring criteria provide a balanced and well-calibrated assessment scale. Notably, quality scores are slightly skewed toward higher values, reflecting that most editing systems tend to preserve audio quality easily, while still facing challenges in precisely executing editing instructions and maintaining content fidelity.
@@
Figure 1b–d further visualize performance across different systems. Notably, Human Editing consistently achieves the highest ratings in all dimensions, establishing an upper bound for automated models. Interestingly, some systems (e.g., sys3, sys5, and sys6) achieve high relevance scores, indicating strong alignment with the editing instructions, yet score lower in faithfulness—highlighting a common trade-off in current models between effective editing and content preservation. This analysis underscores the remaining gap between human and machine editing and suggests a clear direction for future model improvement.
Figure 2: The overall architecture of AuditEval.
Table 2: Performance of the proposed model on the AuditScore dataset at system and utterance levels across three dimensions.
Dimension System-level Utterance-level
MSE↓ LCC↑ SRCC↑ KATU↑ MSE↓ LCC↑ SRCC↑ KATU↑
Quality Score 0.0595 0.6101 0.7004 0.5516 0.1474 0.3318 0.3306 0.2462
Relevance Score 0.0962 0.6196 0.6472 0.4783 0.2963 0.4105 0.3991 0.2886
Faithfulness Score 0.0726 0.8460 0.8809 0.7024 0.2217 0.5327 0.5171 0.3791
Table 3: Comparison of pseudo-parallel data quality before and after filtering. Objective metrics include CLAPt (cross-modal similarity between edited audio and target text), CLAPa(similarity between edited audio and target audio), IS, FAD, FD, and KL divergence (all measuring distributional similarity between edited and original audio). Subjective metrics are predicted human preference scores from AuditEval, used to compare perceived quality and reveal discrepancies with objective measures.
Dataset Objective Subjective (Automatic)
CLAPt(%)↑ CLAPa(%)↑ IS↑ FAD↓ FD↓ KL↓ Quality Relevance Faithfullness
Random1 48.70 82.73 12.23 0.2859 3.207 0.6913 3.365 3.281 3.114
Random2 48.47 82.69 12.23 0.2816 3.197 0.6892 3.365 3.283 3.114
Random3 48.76 82.73 12.20 0.2830 3.205 0.6867 3.366 3.281 3.114
Filtered 49.43 82.76 12.61 0.4587 4.356 0.6783 3.626 3.694 3.493
Method
AuditEval: An Automatic Evaluation Model
Based on our newly curated AuditScore dataset, we develop AuditEval—an automatic MOS-style evaluation model for audio editing built upon the MusicEval framework. AuditEval processes two groups of inputs: the original audio paired with its edited version, and the corresponding textual descriptions of the audio before and after editing, respectively. It outputs three scalar scores reflecting the key evaluation dimensions: Quality, Relevance, and Faithfulness. To account for the different learning complexities across these dimensions, AuditEval employs three independently trained decoding heads, each specialized for predicting one specific score. This simple yet effective design represents a minimal adaptation of the MusicEval framework (see Figure 2), with modest architectural changes and some optimization efforts, while still achieving balanced and reliable predictions across all evaluation aspects.
Pseudo-parallel Data Filtering
@@
Due to the lack of well-established supervised datasets for audio editing, we construct a large-scale (500,000 pairs, 2700h) Pseudo-parallel Audio Editing Dataset following the methodologies of AUDIT (Wang et al. 2023b) and InstructME (Han et al. 2023), and further enhance it using our automatic scoring model, AuditEval. The dataset consists of synthetic audio tuples (A, B, C, A+B, A+C), where Audio A is sampled from clustered AudioCaps (Kim et al. 2019) descriptions, and B/C are single-event clips paired with semantically similar prompts. These are reorganized into instruction-based triplets for Addition, Deletion, and Replacement tasks. A rule-based strategy is used to mix the audio and construct corresponding textual instructions, aiming to maintain acoustic coherence and semantic plausibility. Details are provided in AppendixB C.
@@
However, this kind of rule-based concatenation introduces several limitations: (1) the quality and prompt relevance of individual audio elements are inconsistent; (2) the insertion positions in the background audio are often suboptimal—even when attempting to insert at quieter regions, the resulting mixtures can sound unnatural and make it difficult to distinguish the editing change; (3) the randomly selected A/B/C combinations can be semantically implausible—for example, inserting the sound of paper rustling into a battlefield recording, or replacing male speech with a toilet flushing sound, etc. To address these issues, we apply AuditEval to automatically evaluate and filter synthesized audio editing pairs based on scoring criteria informed by AuditScore. Because AuditScore covers almost all mainstream editing operations including mixing, this filtering is expected to improve data quality and semantic plausibility, advancing toward a practical and scalable supervised dataset (see AppendixC D).
Experiments
Dataset and Metrics
For the automatic evaluation of audio editing, we employ our proposed AuditEval model and train it on the curated AuditScore dataset using a fixed 9:1 train-test split. The dataset will be publicly released to facilitate reproducibility. We report the Mean Squared Error (MSE) at both the utterance and system levels to quantify prediction accuracy. To evaluate consistency with human judgments, we additionally report the Linear Correlation Coefficient (LCC), Spearman’s Rank Correlation Coefficient (SRCC), and Kendall’s Tau Correlation Coefficient (KTAU)(Cooper et al. 2022), which capture linear and rank-based relationships between predicted scores and ground-truth annotations. For the task of selecting high-quality pseudo-parallel audio editing data, we conduct exploratory experiments with objective evaluations to compare the audio quality before and after filtering. Referring to typical training data scales, we retain the top 15% of data for comparison. Specifically, we adopt FAD, FD, and KL, following the methodology of (Liu et al. 2023), to assess the distributional similarity between audio sets. Additionally, we use the CLAP Score to evaluate cross-modal alignment between audio and corresponding descriptions.
(a) Quality
(b) Relevance
(c) Faithfulness
Figure 3: System-level comparison between AuditEval predictions and human ratings across three evaluation aspects.
Results and Analysis
We first evaluate the effectiveness of our automatic evaluation model, AuditEval, which is trained on the curated AuditScore dataset. Table 2 presents the evaluation results at both the system and utterance levels.
At the system level, the mean squared error (MSE) between AuditEval predictions and human ratings remains impressively low—below 0.1 on a 1–5 scale—across all evaluation dimensions. Meanwhile, the linear correlation coefficient (LCC) exceeds 0.6 in every case. Notably, Faithfulness achieves an LCC of 0.84, indicating strong agreement with human judgments. As illustrated in Figure 3, AuditEval effectively ranks different models according to their performance, particularly in terms of executing editing intentions (Relevance) and preserving unedited content (Faithfulness). These results demonstrate that AuditEval successfully fulfills its goal of bridging the evaluation gap in audio editing by providing a scalable and repeatable MOS-style assessment tool to support model comparison and benchmarking.
At the utterance level, the mean squared error (MSE) between AuditEval predictions and human ratings remains below 0.3 (on a 1–5 scale) across all dimensions, with the linear correlation coefficient (LCC) exceeding 0.3 in each case, indicating positive correlation (see Figure 3). Compared to system-level evaluation, utterance-level prediction is inherently more challenging due to increased variability. Nevertheless, AuditEval achieves relatively low MSE values and maintains a reasonable correlation with human judgments, demonstrating its ability to approximate human perception at finer granularity. Importantly, the closeness of AuditEval’s predictions to human ratings in absolute terms lays a solid foundation for downstream applications such as automatic utterance-level filtering or selection.
Among the three dimensions, Faithfulness consistently achieves the highest performance, both at the system and utterance levels. This indicates that the model is relatively effective in evaluating whether the unedited content is preserved after editing. In contrast, Relevance and Quality are more challenging to predict accurately, as they require a deeper joint understanding of both the original and edited audio-text pairs. Unlike standard audio generation evaluation—which typically involves assessing a single audio clip against its corresponding text—audio editing evaluation introduces an additional layer of semantic and acoustic reasoning across paired inputs.
To further assess the effectiveness of our proposed filtering strategy for the Pseudo-parallel Audio Editing Dataset, we compare the quality of three randomly sampled subsets (Random1-3) against a curated subset (Filtered) using standard objective metrics. As shown in Table 3, the Filtered subset outperforms three random baselines across most of the metrics, achieving the highest CLAP-based semantic alignment (CLAP_t: 49.43%, CLAP_a: 82.76%), as well as superior IS (12.61) and KL (0.6783) scores.
(a) Quality
(b) Relevance
(c) Faithfulness
Figure 4: Utterance-level fit between AuditEval predictions and human ratings cross three evaluation aspects.
These results demonstrate that the proposed filtering process significantly improves semantic consistency and editing quality in the pseudo-parallel data.
Nevertheless, certain objective quality metrics, such as FAD and FD, do not fully capture the improvements brought by our filtering strategy. Interestingly, the Filtered subset exhibits noticeably worse FAD and FD scores compared to the random subsets. This discrepancy underscores a fundamental divergence in the focus of subjective and objective evaluations: while human judgments are more attuned to high-level semantic attributes—such as relevance, coherence, and editing accuracy—objective metrics like FAD and FD measure more low-level acoustic properties and may even surpass human perception in such fine-grained distinctions. This distinction explains why semantically grounded metrics (e.g., CLAP, IS) align more closely with human evaluations, whereas acoustic metrics tend to diverge. Therefore, subjective evaluation remains a vital and complementary tool for assessing editing quality, particularly in semantically rich tasks.
Conclusion
In this paper, we address the longstanding challenge of evaluating audio editing systems by introducing AuditEval, a scalable framework for automatic MOS-style assessment. Through the construction of the AuditScore benchmark and the development of the AuditEval model, we enable accurate and efficient quality estimation across multiple dimensions. Furthermore, we demonstrate the practical utility of AuditEval by applying it to curate a high-quality pseudo-parallel dataset, which effectively leads to tangible improvements in data quality. Overall, our work lays a solid foundation for future research in both evaluation and data construction for audio editing.
Appendix A Appendix
Appendix B A. Details of Scoring Criteria Specification

### v2

@@ 3.1 Basic Information of AuditScore
Table 1: Subjective evaluation results across various systems. Detailed distribution of different editing types is presented in Appendix I.
@@
AuditScore is a human-annotated dataset comprising 6,360 pairs of original and edited audio samples, totaling 35.3 hours in duration. These samples are generated using 23 distinct system configurations across seven representative audio editing approaches: SDEdit, DDIM, Edit-Friendly DDPM, CycleDiffusion, AP-adapter, MusicGen, and Human Edit(manual editing). For diffusion-based methods (SDEdit, DDIM, Edit-Friendly DDPM, and CycleDiffusion), we adopt two state-of-the-art generative backbones Tango2  Majumder et al. (2024) and Auffusion  Xue et al. (2024) to synthesize the edited audio. To further promote diversity and enhance model generalization, each configuration is instantiated with three inversion timesteps (75, 125, and 175 out of 200), allowing finer control over the degree of editing. This setup is specifically applied to SDEdit, Edit-Friendly DDPM, and CycleDiffusion to explore a wider range of editing intensities and semantic variations.
@@
To construct a representative and diverse dataset, we first select original audio samples and corresponding captions from AudioCaps. The captions are clustered into 100 semantic groups using GTE-large Li et al. (2023) text embeddings, from which 2 to 3 audio samples are manually selected per cluster. For each selected audio, we design three types of editing instructions, including Addition, Deletion, and Replacement, resulting in a shared pool of 240 instructions across all systems. In addition, each representative systems are assigned 30 unique instructions to further increase the diversity of editing scenarios. Overall, AuditScore covers more than 200 distinct audio event types, with editing operations introducing over 150 novel events. The diversity in algorithms, model architectures, editing parameters, and audio content enables AuditScore to capture a wide range of editing behaviors and quality variations, making it a robust source of expert-informed knowledge for training and evaluating automatic audio editing evaluation models.
@@
To assess the edited audio samples generated by these 23 systems, we conducted a large-scale subjective listening study involving five domain experts with extensive experience in audio dataset curation and perceptual evaluation. Each expert independently rated the editing audio pairs along three key perceptual dimensions: Quality, Relevance, and Faithfulness. The final rating for each utterance was computed as the average of five expert scores across individual dimensions. A comprehensive summary of system IDs and their evaluation results is presented in Table 1.
Figure 1: Distributions of AuditScore annotations (left) and AuditEval predictions (right) at both the system and utterance levels.
3.2 Scoring Protocol and Analysis
@@
To comprehensively evaluate audio editing performance, we adopt a three-dimensional scoring protocol encompassing Quality, Relevance, and Faithfulness. These dimensions are informed by established perceptual evaluation methodologies, such as MOS (Mean Opinion Score), CMOS (Comparison MOS) Loizou (2011), and semantic alignment assessments, and are further tailored to capture the specific nuances of audio editing. Detailed annotation guidelines were developed to ensure consistency across expert raters while preserving flexibility for nuanced perceptual judgments. The scoring scales are as follows(details can be found at Appendix A):
@@
Figure 1 the leftmost column, shows the distribution of scores across the three dimensions. The scores for Quality, Relevance, and Faithfulness each approximately follow a normal distribution, indicating that our scoring criteria provide a balanced and well-calibrated assessment scale. Notably, quality scores are slightly skewed toward higher values, reflecting that most editing systems tend to maintain audio quality easily, while still facing challenges in precisely executing editing instructions and preserving unedited content.
@@
Figure 1 the middle-left column, further visualizes performance across different systems. Notably, Human Editing consistently achieves the highest ratings in all dimensions, establishing an upper bound for automated models. Interestingly, some systems (e.g., sys3, sys5, and sys6) achieve high relevance scores, indicating strong alignment with the editing instructions, yet score lower in faithfulness, highlighting a common trade-off in current models between effective editing and content preservation. This analysis underscores the remaining gap between human and machine editing and suggests a clear direction for future model improvement.
4 Method
Figure 2: Model designs of AuditEval-llm and AuditEval-ssl, along with the corresponding training modules and inference workflows.
Dimensions Models System-level Prediction Utterance-level Prediction
[…]
Quality Obj-Metric 8.7086 0.2277 0.2470 0.1858 - - - -
AuditEval-ssl 0.0595 0.6101 0.7004 0.5516 0.1474 0.3318 0.3306 0.2462
AuditEval-llm 0.0198 0.5814 0.5523 0.3897 0.0999 0.3840 0.3967 0.2922
Relevance Obj-Metric 8.7884 0.2141 0.1779 0.1225 9.0698 0.1344 0.0812 0.0574
AuditEval-ssl 0.0962 0.6196 0.6472 0.4783 0.2963 0.4105 0.3991 0.2886
AuditEval-llm 0.0328 0.7620 0.8142 0.6443 0.2189 0.4766 0.4049 0.2929
Faithfulness Obj-Metric 6.7973 0.8790 0.8843 0.7024 6.9466 0.4981 0.5038 0.3611
AuditEval-ssl 0.0726 0.8460 0.8809 0.7024 0.2217 0.5327 0.5171 0.3791
AuditEval-llm 0.0285 0.9133 0.8952 0.7579 0.1313 0.7213 0.6898 0.5300
Table 2: Performance comparison of score prediction at both system and utterance levels. denotes the Inception Score (IS), distribution-level metric for Quality, denotes the cross-modal CLAP score for Relevance, and denotes the audio–audio CLAP score for Faithfulness.
We present AuditEval, a scalable automatic MOS-style evaluation framework for audio editing that explores two complementary paradigms: an SSL-based predictor (AuditEval-ssl) and an LLM-based evaluator (AuditEval-llm), both producing aspect-wise scores for Quality, Relevance, and Faithfulness.
4.1 AuditEval-SSL
AuditEval-ssl is designed to automatically predict MOS-style scores for audio editing results by leveraging self-supervised audio representations. Given an original audio, an edited audio, and their corresponding original and target descriptions, the model first extracts high-level acoustic features and temporal cues from the audio using a pretrained SSL backbone (we adopt ms-clap, which can automatically align audio and text semantics), and maps these representations to subjective scores along three evaluation dimensions. The core idea of AuditEval-ssl is to transfer the general-purpose audio understanding capability of SSL models to the audio editing evaluation setting, enabling robust and scalable scoring without relying on paired ground-truth edited audio.
4.2 AuditEval-LLM
AuditEval-llm explores an alternative evaluation paradigm in which a large language model serves as an evaluator by reasoning over multimodal evidence. Instead of directly regressing scores from latent embeddings, we convert the input audios together with their textual captions into a structured textual evidence set (details can be found at Appendix H) and prompt the LLM to produce MOS-style ratings. In terms of model design, we adopt the advanced Large Audio Language Model Qwen2.5-Omni-7B. Since the target scores are continuous float values obtained by averaging five independent human ratings, we move beyond conventional dialog-style discrete scoring in a {1–5} set. Specifically, we retain the Qwen2.5-Omni Thinker module and design three post regression heads to predict continuous scores for the three evaluation dimensions. This design leverages the instruction-following, cross-modal understanding, and judgment reasoning capabilities of LLMs for audio editing evaluation.
4.3 Pseudo-parallel Data Filtering
@@
Due to the lack of well-established supervised datasets for audio editing, we construct a large-scalePseudo-parallel Audio Editing Dataset following the methodologies of AUDIT Wang et al. (2023) and InstructME Han et al. (2023), and further enhance it using our automatic scoring model, AuditEval. The dataset consists of synthetic audio tuples (A, B, C, A+B, A+C), where Audio A is sampled from clustered AudioCaps Kim et al. (2019) descriptions, and B/C are single-event clips paired with semantically similar prompts. These are reorganized into instruction-based triplets for Addition, Deletion, and Replacement tasks. A rule-based strategy is used to mix the audio and construct corresponding textual instructions, aiming to maintain acoustic coherence and semantic plausibility. Details are provided in Appendix B.
@@
However, this kind of rule-based concatenation introduces several limitations: (1) the quality and prompt relevance of individual audio elements are inconsistent; (2) the insertion positions in the background audio are often suboptimal—even when attempting to insert at quieter regions, the resulting mixtures can sound unnatural and make it difficult to distinguish the editing change; (3) the randomly selected A/B/C combinations can be semantically implausible—for example, inserting the sound of paper rustling into a battlefield recording, or replacing male speech with a toilet flushing sound, etc. To address these issues, we apply AuditEval to automatically evaluate and filter synthesized editing pairs based on scoring criteria informed by AuditScore. This filtering is expected to improve data quality and semantic plausibility, advancing toward a practical and scalable supervised audio editing dataset (detailed filtering configurations provided in Appendix C).
5 Experiments
Sub-set Filtering Strategy Rule-Based Objective Metrics Human-Aligned Automatic Scores
[…]
Random1 48.70 12.23 0.2859 3.207 0.6913 62.72 5.770 3.601 3.664 4.965
Random2 48.47 12.23 0.2816 3.197 0.6892 62.67 5.769 3.598 3.664 4.964
Random3 48.76 12.20 0.2830 3.205 0.6867 62.74 5.769 3.601 3.667 4.965
Filtered (AuditEval-ssl) 49.43 12.61 0.4587 4.356 0.6783 64.51 5.835 3.619 3.775 5.105
Filtered (AuditEval-llm) 50.96 11.86 0.6003 8.596 0.5289 64.98 5.802 3.771 3.815 5.051
Table 3: Comparison of pseudo-parallel data quality under different filtering strategies across objective metrics and human-aligned scores.
5.1 Experiment Setup
For MOS prediction, we train the proposed AuditEval family of models on the curated AuditScore dataset using a fixed 9:1 train–test split, with detailed training configurations provided in Appendix H. To evaluate consistency with human judgments, we report Mean Squared Error (MSE), Linear Correlation Coefficient (LCC), Spearman’s Rank Correlation Coefficient (SRCC), and Kendall’s Tau Correlation Coefficient (KTAU) at both the utterance and system levels, measuring prediction accuracy as well as rank-based agreement between predicted and ground-truth scores.
For the task of selecting high-quality pseudo-parallel audio editing data, we further conduct exploratory experiments combining objective and automatic subjective evaluations to compare editing quality before and after model-based filtering. In this setting, we retain the top 15% of 59400 samples (over 3,300 hours) according to the predicted scores for comparison. This ratio is chosen as the minimal proportion at which the three random subsets exhibit comparable quality.
Regarding evaluation metrics, following practice in TTA, we adopt IS, FAD, FD, and KL divergence to assess similarity between audios Liu et al. (2023) before and after editing. For objective cross-modal evaluation, we use CLAP Score to measure alignment between edited audio and the corresponding target descriptions. For automatic subjective evaluation, we employ Human-CLAP Takano et al. (2025), which aligns CLAP-based scores with human preferences, as well as AudioBox to provide human-aligned aesthetic quality assessments of edited audio from four perceptual dimensions: Production Quality (PQ), Content Enjoyment (CE), Content Usefulness (CU), and Production Complexity (PC).
5.2 Results and Analysis
Effectiveness of AuditEval.
As shown in Table 2 and the right columns of Fig. 1, both AuditEval-ssl and AuditEval-llm achieve practical and competitive performance in predicting audio editing quality. At the system level, which is commonly used to assess overall model performance and rank different systems, AuditEval-ssl attains a high LCC of around 0.6 on the Quality dimension, together with a low MSE below 0.1 on the 1–5 rating scale. Meanwhile, AuditEval-llm achieves remarkably high correlations of 0.76 and 0.91 on the Relevance and Faithfulness dimensions, respectively, while maintaining MSE values below 0.1, indicating substantially stronger alignment with human perception and clearly outperforming conventional objective metrics.
Compared with system-level evaluation, utterance-level prediction is inherently more challenging due to increased variability and noise at the individual sample level. Nevertheless, across all dimensions, both AuditEval-ssl and AuditEval-llm maintain relatively low MSE values (approximately 0.2 on the 1–5 scale) and preserve reasonable positive correlations (ranging from 0.38 to 0.72) with human judgments. In particular, AuditEval-llm achieves correlations of 0.47 for Relevance and 0.72 for Faithfulness, excelling in semantic related dimensions. These results lay a solid foundation at a finer granularity for subsequent downstream utterance-level data filtering and selection.
AuditEval-ssl vs. AuditEval-llm.
A comparison between AuditEval-ssl and AuditEval-llm (Table 2 and Fig. 1 right) shows that AuditEval-llm achieves state-of-the-art performance across most dimensions and metrics. In particular, on the Relevance and Faithfulness dimensions, which are closely related to semantic understanding, AuditEval-llm demonstrates notably stronger judgment capability. This advantage is likely attributable to the reasoning abilities and intelligence of large audio language models. From an acoustics-oriented perspective, AuditEval-ssl exhibits comparable, and in some metrics even superior, prediction performance on Quality dimension. This aligns with the general observation that audio-specific self-supervised learning models are particularly effective at extracting fine-grained acoustic representations, especially high-frequency cues that influence audio quality.
Effectiveness of Filtering.
As shown in Table 3, the filtered data, selected using either AuditEval-ssl or AuditEval-llm, consistently exhibits improved performance across all human-aligned scores, as well as on CLAP_Score, which is currently one of the most reliable objective metrics for audio editing. Notably, the subjective evaluation models used in our experiments, namely Human-CLAP and AudioBox, are entirely independent of AuditEval in terms of both data and model design. This fully in-the-wild evaluation setting provides strong evidence that the proposed filtering strategy leads to consistent and reliable quality improvements.
Objective Metrics vs. Human-aligned Scores.
As shown in Table 3, improvements in human-aligned scores are not consistently observed in objective metrics, contrary to expectations. Instead, metrics such as FAD and FD, which are commonly used to measure similarity between original and edited audios (faithfulness), exhibit noticeable degradation after filtering. This discrepancy reflects a fundamental difference between subjective and objective evaluation criteria. Human judgments tend to emphasize overall perceptual quality, including coherence, enjoyment, usability, and the semantic correctness of editing operations. In contrast, objective metrics such as FAD and FD primarily focus on low-level acoustic signal variations that are often perceptually insignificant to human listeners. These observations highlight the limitations of objective metrics and underscore the advantages of human-aligned evaluation, particularly for the complex and semantically rich audio editing scenarios.
6 Conclusion
In this paper, we address the long-standing challenges of data scarcity and unreliable evaluation in audio editing by grounding both evaluation and data construction in human judgments. We systematic construct AuditScore, the first human-annotated dataset for MOS-style audio editing assessment, then develop AuditEval-ssl and AuditEval-llm to study two different paradigms for automatic subjective evaluation and high-quality dataset mining, providing both practical tools and methodological insights. Through comprehensive experiments, we reveal key challenges in audio editing evaluation and data curation, and highlight the role of human-aligned evaluation in improving evaluation and data quality.
Appendix A Details of Scoring Criteria Specification